OpenStudy (vera_ewing):

The temperature of a chemical reaction ranges between 30°C and 70°C. The temperature is at its lowest point when t = 0 and completes one cycle over a six-hour period. What is a sine function that would model this reaction?

OpenStudy (vera_ewing):

@Michele_Laino f(t) = 50 sin 10t + 20 ?

OpenStudy (michele_laino):

the amplitude of your function, is: \[\Large A = \frac{{\max value - \min value}}{2} = \frac{{70 - 30}}{2} = ...?\]

OpenStudy (vera_ewing):

40/2 = 20

OpenStudy (vera_ewing):

f(t) = 20 sin 10t + 50 is the answer?

OpenStudy (michele_laino):

we can write: \[\Large f\left( t \right) = 20\sin \left( {Bt} \right) + 50\] since we have these limit values: \[\Large 30 \leqslant 20\sin \left( {Bt} \right) + 50 \leqslant 70\]

OpenStudy (michele_laino):

now we have to determine the value of the constant B

OpenStudy (vera_ewing):

So it's f(t) = 20 sin pi/5 t + 50 ?

OpenStudy (michele_laino):

I think that we have: \[\Large B = \frac{{2\pi }}{6}\] since our function has to assume the same values after 6 hours

OpenStudy (michele_laino):

therefore: \[\Large f\left( t \right) = 20\sin \left( {\frac{{2\pi }}{6}t} \right) + 50\] is the right function

OpenStudy (vera_ewing):

Those are my answer choices ^

OpenStudy (michele_laino):

please wait a moment, I'm pondering...

OpenStudy (michele_laino):

please look at this drawing:

OpenStudy (michele_laino):

maybe there is a typo since we can rewrite our function as below: \[\Large f\left( t \right) = 20\sin \left( {\frac{\pi }{3}t} \right) + 50\]

OpenStudy (michele_laino):

please wait a moment, I understand

OpenStudy (michele_laino):

I confirm my answer above

OpenStudy (vera_ewing):

So which one should I choose?

OpenStudy (michele_laino):

try with the last option, please

OpenStudy (michele_laino):

is the period T=10 hours?

OpenStudy (michele_laino):

I think that there is a typo, namely: T=10 hours and not T=6 hours